1. Customers who are registered on a corporate web site are summarized by the type of shipping contract they use and the number of orders in the previous month. The number of customers in each category are shown in the following table. Shipping Contract No orders One order More than one order 40 Total Express 25 65 Standard 15 45 20 80 Total 40 110 60 210 130 (a) If customers are selected at random without replacement, how many unique samples of eight express and ten standard customers are possible? Order does not matter. (2 pts) (b) Suppose 10 express and 15 standard customers are selected without replacement. What is the probability of selecting exactly one express customer with no orders last month and exactly one standard customer with no orders last month? (3 pts) (c) Suppose that three customers are selected randomly, without replacement, for a survey. What is the probability that at least two selected customers use express shipping? (4pts) (d) What is the probability that a randomly selected customer uses standard shipping or had no orders? (4 pts)
Answer #1

a)Given that the customers are selected at random, the number of unique samples of selecting eight express and ten standard customers are ,

The number of favourable cases of selecting 8 express customers out of 130 express customers is 130C8

The number of favourable cases of selecting 10 standard customers out of 80 express customers is 80C10

The number of unique samples of selecting eight express and ten standard customers are = 130C8 * 80C10

b)Given that 10 express and 15 standard customers are selected without replacement, the probability of selecting one express customer with no orders last month and exactly one standard customer with no orders last month is

Number of express customer with no orders last month is 25

Number of express customer with no orders last month is 15

exhaustive number of cases = 210C25 (Since 25 customers are to be selected from 210 customers)

favourable number of cases = 25C1* 105C9 * 15C1 * 65C14

Required probability = favourable number of cases / exhaustive number of cases

                              = 25C1* 105C9 * 15C1 * 65C14 / 210C25 = 0.000972

From 10 express and 15 standard customers selected without replacement, the probability of selecting one express customer with no orders last month and exactly one standard customer with no orders last month is 0.000972                 

c)Given that three customers are selected randomly without replacement, the probability that at least two selected customers uses express shipping is,

P(at least 2 selected customers uses express shipping)=P( 2 or more selected customers uses express shipping)

=P( 2 selected customers uses express shipping) +P( 3 selected customers uses express shipping)

130C2*80 C1130Ca 210 C 210C

130 129 80 30 129-128 210 209 208 3 2 1 3 241

3577600,440875 0.440875 0.235133 0.676008 1521520 1521520

From the three customers selected randomly without replacement, the probability that at least two selected customers uses express shipping is 0.676008

d) The probability that a randomly selected customer uses standard shipping or had no orders

A - event of randomly selecting a customer who uses standard shipping

8o P(A) A) P(A) = n(S)-210

Where n(A) is the favourable number of cases and n(S) is the exhaustive number of cases

B - event of randomly selecting a customer who had no orders

P(B) = n(B) 5 n(S) 210

The probability that a randomly selected customer uses standard shipping or had no orders is P(Aor B) = P(A U B)

P(AU B) = P(A) + P(B) _ P(An B)

The probability that a randomly selected customer uses standard shipping and had no orders isP(A and B)=P(Acap B)

n(Pn B 15 P(An B) = n(PNB) 15 _ n(S 210

P(AU B) = P(A) + P(B) _ P(An B)

105 80 40 15 210 210210 210 P(AUB)

The probability that a randomly selected customer uses standard shipping or had no orders is 0.5.